# A Survey of Deep Learning for Complex Speech Spectrograms

### v1

@@ Abstract
Recent advancements in deep learning have significantly impacted the field of speech signal processing, particularly in the analysis and manipulation of complex spectrograms. This survey provides a comprehensive overview of the state-of-the-art techniques leveraging deep neural networks for processing complex spectrograms, which encapsulate both magnitude and phase information. We begin by introducing complex spectrograms and their associated features for various speech processing tasks. Next, we explore the key components and architectures of complex-valued neural networks, which are specifically designed to handle complex-valued data and have been applied for complex spectrogram processing. We then discuss various training strategies and loss functions tailored for training neural networks to process and model complex spectrograms. The survey further examines key applications, including phase retrieval, speech enhancement, and speech separation, where deep learning has achieved significant progress by leveraging complex spectrograms or their derived feature representations. Additionally, we examine the intersection of complex spectrograms with generative models. This survey aims to serve as a valuable resource for researchers and practitioners in the field of speech signal processing and complex-valued neural networks.
@@
Complex-valued neural networks , deep learning , complex spectrogram processing , phase retrieval , speech enhancement , speech separation , generative models
journal: Astronomy &\&& Computing
@@ 1 Introduction
Figure 1: This figure shows the survey’s structure, organized around complex spectrogram modeling and processing. The framework is divided into four main components: (1) Complex-Valued Neural Networks, with a focus on those specifically designed to process and analyze complex-valued spectrograms; (2) Training Strategies and Loss Functions, which analyze optimization techniques and loss function designs tailored for learning complex spectrogram representations; (3) Applications, highlighting the practical utility of complex spectrogram research across various domains, including phase retrieval, speech enhancement, and speech separation; (4) Generative Models, covering VAEs, GANs, and diffusion models in the context of complex spectrogram-based tasks. Notably, components (2), (3), and (4) include works based on both real-valued and complex-valued neural networks.
@@
Among these representations, early studies predominantly focused on the magnitude spectrogram, for several key reasons. First, magnitude spectrograms exhibit well-defined structural properties, which facilitate effective feature extraction in deep neural network (DNN)-based approaches. Second, magnitude spectrograms have played a significant role in conventional speech signal processing, leading mordern DNN-based approaches to naturally inherit this focus. In contrast, phase spectrograms have received comparatively less attention, primarily due to the challenges associated with their modeling. These difficulties stem from the inherent properties of phase spectrograms, including the lack of well-defined structure, the issue of phase wrapping, and a high sensitivity to temporal shifts in the waveform. Besides, there has been skepticism regarding the significance of phase spectrograms. Early studies, such as Ephraim and Malah (1984) and Wang and Lim (1982), argued that phase spectrograms are not essential, thereby supporting on magnitude-only speech processing approaches. However, subsequent research has led to a reassessment of this view. For example, Paliwal et al. (2011) demonstrated that enhancing phase information can improve perceptual quality, even when the corresponding magnitude spectrogram is corrupted. Moreover, Le Roux (2011) showed that when a magnitude spectrogram and an irrelevant phase spectrogram are transformed back to the time domain, the resulting audio waveform can significantly exceed expectations. These findings underscore the critical role of phase information in speech signal processing and have spurred growing interest in incorporating phase modeling into modern approaches.
Consequently, recent research has increasingly focused on processing complex spectrograms using DNN-based methods. Given that complex spectrograms can be decomposed into either magnitude–phase or real–imaginary representations, a variety of approaches have been proposed. Magnitude–phase spectrograms have been widely studied, particularly in tasks such as phase retrieval and speech enhancement. Among existing methods, real-valued neural networks remain the dominant implementation choice for processing magnitude–phase representations. Furthermore, due to the inherent difficulty of accurately estimating phase information (Williamson et al. (2016)), the real–imaginary representation has gained popularity and is now widely adopted across a broad range of applications. Both real-valued and complex-valued neural networks have been explored for handling real-imaginary representations. When using real-valued networks, the real and imaginary components are typically either concatenated into a single input vector or treated as two separate input channels. In contrast, complex-valued neural networks process all elements — including inputs, outputs, parameters, and internal operations — in the complex domain.
Research on complex-valued neural networks can be traced back to Widrow et al. (1975), where the least mean square algorithm was extended to the complex-valued domain. Over the past few decades, research on complex-valued neural networks has been ongoing, and interest in this topic has grown. Complex-valued neural networks have been applied widely, including but not limited to, audio and speech signal processing, image processing (Quan et al. (2021); Rawat et al. (2021)), radar signal processing (Yu et al. (2019); Fuchs et al. (2021)), object discovery (Löwe et al. (2022)) and so on.
Several studies have theoretically demonstrated the potential advantages of complex-valued neural networks from multiple perspectives. Hirose (2011) pointed out that a complex-valued neural network has a solution space with a lower degree of freedom. In addition, Voigtlaender (2023) extended the universal approximation theory to complex-valued domain, and proved that with a proper activation function, both shallow and deep complex-valued neural networks can uniformly approximate any continuous function. Moreover, Nitta (2002, 2013) illustrated the advantage of complex-valued neural networks at critical points and local minima. Recent research, such as (Tan et al. (2022); Wu et al. (2024)), continues to explore and deepen our understanding of complex-valued neural networks. Even though the aforementioned studies highlight the theoretical advantages of complex-valued neural networks over their real-valued counterparts, a definitive conclusion regarding their comparative performance remains elusive.
Recent studies in speech signal processing are increasingly focusing on using real-valued neural networks to handle complex spectrogram processing, especially when using discriminative approaches. This may be because real-valued neural networks have less computational cost and can achieve comparable performance. Wu et al. (2023) made adequent experiments to compare the performance between the widely-used complex-valued backbone DCCRN with its real-valued counterpart. Results demonstrate that complex-valued neural networks slightly outperform the real-valued ones, but with greater computational cost.
Nevertheless, we believe that it is still too early to draw a definitive conclusion regarding the performance of real-valued versus complex-valued neural networks in handling complex spectrograms. First of all, research on complex-valued neural networks is ongoing, and our understanding of these networks continues to deepen. Second, complex-valued neural networks have been instrumental in recent advancements applying generative models (such as diffusion models and variational autoencoders) to complex spectrograms.
Thus, in this survey, we provide a comprehensive review of how deep learning approaches address issues related to complex spectrograms. The structure of this survey is divided into two main parts. First, we review the most widely used complex-valued neural networks in speech signal processing. Then, we examine how deep learning approaches are applied to tackle problems associated with complex spectrograms. It is important to emphasize that the second part not only includes studies on complex-valued neural networks but also covers work utilizing real-valued neural networks to process complex spectrograms, as the two approaches complement and advance each other.
We aim for this work to contribute into the following aspects:
1. Complex-valued neural network
The complex spectrogram, as a fundamental analytical feature of speech signals, represents a promising application area for complex-valued neural networks. This survey provides a comprehensive overview of the most widely used complex-valued neural networks in speech signal processing, covering key components and the architectures. We intend for this to be a valuable resource for researchers exploring complex-valued neural networks.
2. Complex spectrogram processing
In this survey, we systematically review and analyze recent advancements in training strategies and loss functions for learning complex spectrograms. In addition, we examine three key application areas — phase retrieval, speech enhancement, and speech separation — within the framework of complex spectrogram processing. We further discuss emerging approaches that leverage generative models for processing complex spectrograms. This work aims to provide valuable insight and serve as a guide for future research in the field of complex spectrogram processing.
@@ 2. Complex spectrogram processing
The organization of this paper is illustrated in Figure 1. Section 2 introduces the representation of complex spectrograms and their associated features. Section 3 outlines the key components of complex-valued neural networks. Section 4 reviews representative architectures of complex-valued neural networks, including convolutional neural networks, recurrent neural networks, and transformers, all of which have been applied to various tasks in speech signal processing. Section 5 explores training strategies and loss functions specifically designed for learning complex spectrogram representations. Sections 6 provide a comprehensive survey of state-of-the-art methods for phase retrieval, speech enhancement, and speech separation based on complex spectrogram processing. Section 7 summarizes recent advancements in generative models operating on complex spectrograms. Finally, Section 8 concludes the paper.
In addition, Table 1 presents the mathematical notations used throughout this paper, while Table 2 outlines the common abbreviations related to speech signals and deep learning referenced herein.
@@ 2 Complex Spectrogram Representation
Figure 2: The figure presents features derived from one utterance in the TIMIT dataset (speaker ID: FVMH0, utterance ID: SA1), including log-magnitude, phase, absolute real/imaginary components, group delay, and instantaneous frequency spectrograms. These time-frequency representations were computed using a Short-Time Fourier Transform (STFT) with a 25 ms Hamming window and 10 ms frame shift. To enhance visualization clarity, both real and imaginary spectrogram components are displayed using their absolute values on the logarithmic scale. Comparative analysis reveals that the log-magnitude spectrogram, complex component representations (real/imaginary), group delay, and IF spectrograms demonstrate significantly clearer structural patterns than the phase spectrogram.
@@
Research working on magnitude-phase spectrograms is also ongoing, but challenges are mostly posed in phase spectrogram processing. This is typically attributed to the fact that phase is wrapped into the range of $[-\pi,\pi]$, and it is hard to find clear structure. Therefore, derivatives of phase are often used to bypass directly dealing with phase itself. The derivatives of phase spectrogram $\theta(l,k)$ over time is called instantaneous frequency (IF):
@@
and over frequency, which is referred to as group delay (GD),
@@ 3 Key Components of Complex-Valued Neural Networks
In this section, we will briefly introduce the complex-valued activation functions and complex-valued batch normalization that appear in complex-valued neural networks when dealing with complex spectrograms.
@@ 3.1.2 $\mathbb{C}$ReLU
3.1.3 $z$ReLU
Another ReLU-based complex activation function was proposed by Guberman (2016) and is referred to as $z$ReLU in Trabelsi et al. (2018). The definition of $z$ReLU is as below:
@@ 4 Complex-Valued Neural Networks
In this section, we will introduce complex-valued neural networks, covering various architectures including feed-forward neural networks, recurrent neural networks, convolutional neural networks, and transformers.
@@ 4.3 Complex-Valued Recurrent Neural Networks
Given that previous studies on complex-valued RNNs have shown positive results, Wolter and Yao (2018) was the first to specifically investigate a complex RNN architecture that operates directly on complex-valued spectrograms. Wolter and Yao (2018) first defined a basic formulation for a complex RNN as:
@@ 4.4 Complex-Valued Transformers
[…]

### v2

@@ Abstract
Recent advancements in deep learning have significantly impacted the field of speech signal processing, particularly in the analysis and manipulation of complex spectrograms. This survey provides a comprehensive overview of the state-of-the-art techniques leveraging deep neural networks for processing complex spectrograms, which encapsulate both magnitude and phase information. We begin by introducing complex spectrograms and their associated features for various speech processing tasks. Next, we examine the key components and architectures of complex-valued neural networks, which are specifically designed to handle complex-valued data and have been applied to complex spectrogram processing. As recent studies have primarily focused on applying real-valued neural networks to complex spectrograms, we revisit these approaches and their architectural designs. We then discuss various training strategies and loss functions tailored for training neural networks to process and model complex spectrograms. The survey further examines key applications, including phase retrieval, speech enhancement, and speaker separation, where deep learning has achieved significant progress by leveraging complex spectrograms or their derived feature representations. Additionally, we examine the intersection of complex spectrograms with generative models. This survey aims to serve as a valuable resource for researchers and practitioners in the field of speech signal processing, deep learning and related fields.
@@
Complex spectrogram processing , deep learning , complex-valued neural networks , low latency , phase retrieval , speech enhancement , speaker separation , generative models
journal: Astronomy &\& Computing
@@ 1 Introduction
Figure 1: This figure illustrates the overall structure of the survey, centered on the modeling and processing of complex spectrograms. The framework is divided into five main components: (1) complex-valued neural networks specifically designed for complex spectrogram processing; (2) RVNNs applied to complex spectrograms; (3) training strategies and loss functions, encompassing low-latency methods, optimization techniques, and loss designs tailored to complex spectrogram learning; (4) applications, covering speech enhancement, speaker separation, and phase retrieval; and (5) generative models, including approaches based on VAEs, GANs, and diffusion models for complex spectrogram processing.
@@
Among these representations, early studies focused predominantly on the magnitude spectrogram for several key reasons. First, magnitude spectrograms exhibit well-defined structural properties and facilitate effective feature extraction in deep neural network (DNN)-based approaches. Second, magnitude spectrograms have played a significant role in conventional speech signal processing, leading mordern DNN-based approaches to naturally inherit this focus. In contrast, phase spectrograms have received comparatively less attention, primarily due to the challenges associated with their modeling. These difficulties stem from the inherent properties of phase spectrograms, including the lack of well-defined structure, the issue of phase wrapping, and a high sensitivity to temporal shifts in the waveform. Besides, there has been skepticism regarding the significance of phase spectrograms. Early studies, such as Ephraim and Malah (1984) and Wang and Lim (1982), argued that phase spectrograms are not essential in speech processing. However, subsequent research has led to a reassessment of this view. For example, Paliwal et al. (2011) demonstrated that enhancing phase information can improve perceptual quality. Moreover, Le Roux (2011) showed that combining a magnitude spectrogram with an irrelevant phase yields an unexpected audio waveform. These findings underscore the critical role of phase information and have spurred growing interest in incorporating phase modeling into modern approaches. Moreover, reconstruction of waveforms from estimated magnitude spectrograms commonly relies on classical algorithms such as Griffin–Lim or on neural vocoders, both of which are prone to introducing additional distortion. This further highlights the limitations of magnitude-only approaches and emphasizes the significance of complex spectrogram processing.
Meanwhile, time-domain signals, preserving all speech information, have been widely adopted in various studies. The main advantages of processing time-domain signal include bypassing the STFT, thereby avoiding explicit phase spectrogram modeling and enabling low-latency processing. Although representative time-domain models like Conv-TasNet (Luo and Mesgarani (2019)) have shown promising performance in speaker separation, recent studies (Zhang et al. (2024b); Wang et al. (2023a)) indicate that time-frequency domain models can achieve higher performance. Furthermore, Wang et al. (2023b) observed that although Conv-TasNet employs an encoder to generate “pseudo-STFT” features, these features are not narrowband, further limiting Conv-TasNet compatibility with conventional STFT-based enhancement methods (e.g., beamforming, weighted prediction error) that rely on narrowband assumptions. In contrast, complex spectrogram–based models integrate seamlessly with such techniques, making them suitable for applications like multi-channel speech enhancement. Moreover, while complex spectrogram–based methods are limited by the time–frequency resolution trade-off of the STFT and typically incur longer delays than time-domain approaches, recent studies have substantially reduced their latency. The latest models achieve delays as low as 4 ms, enabling their deployment in real-time applications (Wang et al. (2023b); Wu and Braun (2025)).
Increasingly, research has focused on DNN-based methods for complex spectrogram processing, implemented with real-valued (RVNNs) or complex-valued neural networks (CVNNs). From the perspective of deep learning development, RVNNs remain the mainstream and have long been widely applied, while CVNNs has also garnered increasing attention in recent years. Research on CVNNs can date back to Widrow et al. (1975). Since then, CVNNs have been applied in diverse areas such as audio processing, image denoising (Quan et al. (2021); Rawat et al. (2021)), radar signal processing (Yu et al. (2019); Fuchs et al. (2021)), and object discovery (Löwe et al. (2022)). Several studies have highlighted theoretical advantages of CVNNs (Tan et al. (2022); Wu et al. (2024); Hirose (2011); Voigtlaender (2023); Nitta (2002, 2013)). Nevertheless, despite established theoretical benefits, the practical superiority of CVNNs over RVNNs remains inconclusive. Recent studies in speech signal processing are mostly focusing on using RVNNs to handle complex spectrogram processing. This may be because RVNNs have lower computational cost while achieving comparable performance (Wu et al. (2023)).
Using deep learning for complex spectrogram processing has yielded promising results across various related fields closely linked to perceptual quality, including phase retrieval, speech enhancement, and speaker separation. Progress has been made in architectural design, training strategies, and loss function development throughout the exploration. Moreover, recent studies have increasingly focused on low-latency processing, universal models, further performance improvements, and lower computational cost, among others. There is also growing interest in leveraging generative models for complex spectrogram processing.
However, despite these advances, the field remains fragmented and a systematic survey is still lacking. This survey aims to fill the gap and provide a comprehensive summary of DNN-based approaches for complex spectrogram processing, serving as a reference for both future research and practical applications.
This work begins by introducing the complex spectrograms and their associated features. It then provides a summary of methods for processing complex spectrograms using CVNNs, primarily from the following perspectives: (1) Research on CVNNs is ongoing, and our understanding of their properties and capabilities continues to evolve, and (2) complex-valued operations possess inherent characteristics, such as the ability to represent data in both Cartesian and polar coordinates, and to simultaneously perform scaling and rotation. Subsequently, this review examines the use of RVNNs for complex spectrogram processing, including input handling and architectural designs. Rather than advocating for a particular network type, we here emphasizes that all conclusions should be grounded in experimental results and practical deployment considerations. This work also investigates the application of deep learning to problems associated with complex spectrograms, encompassing low-latency techniques, optimization strategies, and specialized loss functions. Additionally, we examine three key application areas — phase retrieval, speech enhancement, and speaker separation — within the framework of complex spectrogram processing. Emerging methods leveraging generative models are also summarized. Overall, this work aims to offer a comprehensive foundation for researchers developing DNN-based methods for complex speech spectrograms.
The structure of this survey is shown in Figure 1. In addition, Table 1 presents the mathematical notations used throughout this paper, while Table 2 outlines the common abbreviations related to speech signals and deep learning referenced herein.
@@
CVNN Complex-valued neural network
RVNN Real-valued neural network
@@ 2 Complex Spectrogram Representation
((a)) Log-magnitude
((b)) Phase
((c)) Group Delay
((d)) Real (absolute value on logarithm scale)
((e)) Imaginary (absolute value on logarithm scale)
((f)) Instantaneous Frequency
Figure 2: Features derived from one utterance in the TIMIT dataset (speaker ID: FVMH0, utterance ID: SA1). Time-frequency representations include log-magnitude, phase, real/imaginary components, group delay, and instantaneous frequency spectrograms. Computed using STFT with a 25 ms Hamming window and 10 ms frame shift. For clarity, real and imaginary spectrograms are displayed as absolute values on the logarithmic scale.
@@
Research working on magnitude-phase spectrograms is also ongoing, but challenges are mostly posed in phase spectrogram processing. This is typically attributed to the fact that phase is wrapped into the range of $\pi,\pi]$, and it is hard to find clear structure. Therefore, derivatives of phase are often used to bypass directly dealing with phase itself. The derivatives of phase spectrogram $\theta(l,k)$ over time is called instantaneous frequency (IF; Stark and Paliwal (2008)):
@@
and over frequency, which is referred to as group delay (GD; Yegnanarayana and Murthy (1992)),
@@ 3 Key Components of Complex-Valued Neural Networks
In this section, we will briefly introduce the complex-valued activation functions and complex-valued batch normalization that appear in CVNNs when dealing with complex spectrograms.
@@ 3.1.2 $\mathbb{C}$ReLU
3.1.3 zzReLU
Another ReLU-based complex activation function was proposed by Guberman (2016) and is referred to as zzReLU in Trabelsi et al. (2018). The definition of zzReLU is as below:
@@ 4 Complex-Valued Neural Networks
In this section, we will introduce CVNNs, covering various layer types, including feed-forward neural networks, recurrent neural networks, convolutional neural networks, and transformers. For details on how these layers are integrated into complete architectures (e.g., DCUNET, DCCRN), please refer to Section 7.
@@ 4.3 Complex-Valued Recurrent Neural Networks
Given that previous studies on complex-valued RNNs have shown positive results, Wolter and Yao (2018) were the first to specifically investigate a complex RNN architecture that operates directly on complex-valued spectrograms. Wolter and Yao (2018) first defined a basic formulation for a complex RNN as:
@@ 4.4 Complex-Valued Transformers
$\text{Min-Max-Norm}(\mathbf{Z})=\frac{\mathbf{Z}-\min(\mathbf{Z})}{\max(\mathbf{Z})-\min(\mathbf{Z})}.$ (30)
@@
5 Using Real-Valued Neural Networks for Complex Speech Spectrograms
Compared with CVNNs, RVNNs are more mature and widely studied. In this section, we briefly introduce approaches that employ RVNNs for complex spectrogram processing. Since a complex spectrogram can be represented in either the Cartesian coordinate system (real and imaginary components) or the polar coordinate system (magnitude and phase components), different strategies have been proposed. We begin by examining the inputs to real-valued neural networks, then proceed to the architectural designs for handling the real and imaginary components of spectrograms, and finally discuss architectures that operate on magnitude and phase representations.
5.1 Input Representations: Concatenated and Stacked
((a)) Concatenated
((b)) Stacked
Figure 3: Illustration of two common input representations for real-imaginary or magnitude-phase spectrograms in RVNN architectures. (a) Concatenated representation (Concatenated): The real and imaginary (or magnitude and phase) components of each complex-valued time–frequency vector are concatenated into a single vector, resulting in a higher-dimensional real-valued feature vector. (b) Stacked representation (Stacked): the real and imaginary (or magnitude and phase) components are stacked into a three-dimensional tensor with dimensions [Frequency, Time, 2], where the last dimension corresponds to the real and imaginary (or magnitude and phase) components.
Depending on the architectural design, the front-end of RVNNs typically process the complex spectrograms using one of the two main strategies.
The first strategy is the concatenated vector representation, where the real-imaginary or magnitude-phase components of each complex-valued time–frequency vector are concatenated into a single higher-dimensional vector before being fed into the network, as illustrated in Fig. 3(a). Representative works adopting this approach include Williamson and Wang (2017b), Fan et al. (2024), FullSubNet (Hao et al. (2021)), USR-BSRNN (Yu and Luo (2023)), and MambAttention (Kühne et al. (2025a)).
The second strategy is the stacked representation, as shown in Fig. 3(b). In this case, a complex spectrogram of size $[Frequency,Time]$ is extended to $[Frequency,Time,2]$, where the additional dimension corresponds to the real and imaginary components or the magnitude and phase spectrograms. The stacked tensor is then treated as a multi-channel real-valued input, analogous to the RGB channels of an image. Representative works include Fu et al. (2017), CRN (Tan and Wang (2019b)), CMGAN (Cao et al. (2022)) and MP-SENet (Lu et al. (2023)).
Overall, the concatenated vector representation is commonly used in architectures with a fully connected layer or RNN layer as the front end (i.e., the first layer is fully connected or RNN), whereas the stacked representation arranges the spectrogram components channel-wise — treating the real, imaginary or magnitude, phase parts as separate components — and is therefore particularly suitable for architectures with a convolutional front end. Subsequent layers (e.g., CNN, RNN or attention layers) can be incorporated into the framework to further improve performance.
5.2 Architectural Design
5.2.1 For Real-Imaginary Spectrograms
In early efforts to process real-imaginary spectrograms using RVNNs, fully connected layers were directly applied (Williamson and Wang (2017b)). Fu et al. (2017) were among the first to introduce a CNN framework for processing complex speech signals.
Speech signals exhibit clear temporal characteristics, with strong dependencies between contextual frames, and uneven energy distribution across frequency bands. Therefore, subsequent research has often adopted methods that combine global and local information to enhance model performance. For example, CRN (Tan and Wang (2019b)) employs a U-Net-like encoder-decoder structure with skip connections across the encoder and decoder, and incorporates RNN-based blocks to capture global temporal dependencies. Similar improved models include PoCoNet (Isik et al. (2020)) and DPCRN (Le et al. (2021)).
Moreover, modeling of full-band and sub-band information has also received widespread attention. Full-band models can capture global information but lack dedicated mechanisms for modeling sub-band features, such as signal stationarity and local spectral patterns. Conversely, subband models (Li and Horaud (2020)) are the exact opposite. FullSubNet (Hao et al. (2021)) addresses this issue by sequentially connecting full-band and sub-band models and jointly training them to integrate the advantages of both.
Following the dual-path RNN architecture (Luo et al. (2020)), many subsequent models have leveraged both sub-band and cross-band modeling to more effectively capture temporal and spectral information. In contrast to FullSubNet which sequentially connects full-band and sub-band models for joint learning, dual-path approaches typically integrate a sub-band module and a cross-band module within a single block. This block is then stacked multiple times to progressively refine the representations. Representative models include DPT-FSNet (Dang et al. (2022)), CMGAN (Cao et al. (2022)), TF-GridNet (Wang et al. (2023a)), SpatialNet (Quan and Li (2024)), and TF-CrossNet (Kalkhorani and Wang (2024)), which effectively capture fine-grained local spectral details while simultaneously modeling global time–frequency dependencies.
Besides, the inherent spectral energy imbalance across frequencies can lead to distortion and information loss in high-frequency components. To address this, the Band-Split RNN (BSRNN) was proposed (Luo and Yu (2023)). In this approach, the input signal is split into several sub-bands. A bi-directional LSTM is employed to process information in the lower-frequency bands (<<8 kHz), while a uni-directional LSTM is used to capture high-frequency information (>>8 kHz).
More recently,increasing attention has been directed toward the development of universal models that can robustly process audio signals under diverse conditions, such as varying sampling rates and multi-channel configurations. Notable examples include USR-BSRNN (Yu and Luo (2023)) and Zhang et al. (2023), among others.
5.2.2 For Magnitude-Phase Spectrograms
RVNNs for processing magnitude and phase spectrograms often adopt a dual-branch (also referred to as dual-path [^1] Two distinct ”dual-path” mechanisms have been proposed in the literature. The first is exemplified by the dual-path RNN, which incorporates intra- and inter-chunk processing to capture both local and global information. Representive works include dual-path RNN (Luo et al. (2020)) and TF-GridNet (Wang et al. (2023a)) among others. The second dual-path mechanism, noted in this work as dual-branch, discussed here, employs a parallel architecture, where magnitude and phase information are processed in separate branches. ) architecture, as shown in Fig 4 (a). In this framework, one branch is dedicated to modeling the magnitude spectrogram, while the other focuses on the complex (Fu et al. (2022)) or phase (Yin et al. (2020)) spectrogram. Since phase information is inherently more difficult to model, effective interaction between the two branches is typically required, and multiple information exchanges are often introduced across different layers to facilitate joint learning. Representative works under this line include PHASEN (Yin et al. (2020)), UFormer (Fu et al. (2022)), and DB-AIAT(Yu et al. (2022a)).
Recently, there exists an encoder-decoder-based architecture for processing magnitude-phase spectrograms, as shown in Fig 4 (b). Within this framework, both magnitude and phase spectrograms are first encoded into feature representations by an encoder. These features are then fed into a two-stage block, which processes them sequentially along the temporal and frequency dimensions. The block is typically instantiated with models such as Conformer (Gulati et al. (2020)), Mamba (Gu and Dao (2023)), or xLSTM (Beck et al. (2024)). Finally, the processed features are passed to the magnitude decoder and phase decoder, where the corresponding spectrogram components are reconstructed. Representative works include MPSENet (Lu et al. (2023)) and MambaAttantion (Kühne et al. (2025a)).
((a)) Dual-Branch
((b)) Encoder-Decoder
Figure 4: Illustration of two commonly used architectures for processing magnitude and phase spectrograms.